# Semi-Supervised Deep Sobolev Regression: Estimation, Variable Selection and Beyond

### v1

@@ 5.1 Analysis for Deep Sobolev Regressor
Lemma 5.2 (Oracle inequality).
Suppose Assumptions 1 to 5 hold. Let $\widehat{f}_{{\mathcal{D}}}^{\lambda}$ be the deep Sobolev regressor defined as (11). Then it follows that for each $\lambda>0$ and $n\geq\operatorname{VCdim}({\mathcal{F}})$,
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}$ $\lesssim\beta\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}({\mathcal{F}},n),$
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{]}$ $\lesssim\beta\lambda+\lambda^{-1}\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\lambda^{-1}\varepsilon_{\operatorname{gen}}({\mathcal{F}},n),$
@@
$\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)$ $=\inf_{f\in{\mathcal{F}}}\Big{\{}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2}+\lambda\|\nabla(f-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{\}},$
$\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)$ $=\inf_{\delta>0}\Big{\{}(B_{0}^{2}+\sigma^{2})\frac{\operatorname{VCdim}({\mathcal{F}})}{n\delta}+(B_{0}+\sigma)\delta\Big{\}}.$
As discussed in Section 4, Györfi et al. (2002, Chapter 21) has investigated an optimization problem similar to the deep Sobolev regressor. However, to the best of our knowledge, we are the first to demonstrate the oracle inequality for the gradient of estimator. The proof employs a similar technique as that of Lemma 5.1. Specifically, the deep Sobolev regressor acts as the minimizer of (11), which implies that it satisfies a variational inequality derived from the first-order optimality condition (Hinze et al., 2009; Tröltzsch, 2010). By utilizing standard techniques from statistical learning theory, we are able to derive the desired oracle inequality.
In simple terms, if we select an appropriate neural network class and have a sufficiently large number of labeled data pairs, we can make the approximation error and generalization error arbitrarily small. Consequently, the overall error is primarily determined by the regularization parameter $\lambda$. At this point, the error bound aligns with rates in Lemma 5.1.
Recall the oracle inequality derived in Lemma 4.1, which requires the neural network to approximate the regression function while restricting its gradient norm. In contrast, the approximation term in Lemma 5.2 necessitates the neural network to approximate both the regression function and its derivatives simultaneously. Thus, we now introduce the following approximation error bound in $H^{1}$-norm.
Lemma 5.3 (Approximation in $H^{1}$-norm).
Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. Set the hypothesis class as a deep ReQU neural network ${\mathcal{F}}={\mathcal{N}}(L,S)$ with $L={\mathcal{O}}(\log N)$ and $S={\mathcal{O}}(N^{d})$. Then for each $\phi\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 2}$, there exists $f\in{\mathcal{F}}$ such that
@@
Theorem 5.4 (Convergence rates).
Suppose Assumptions 1 to 5 hold. Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. Assume that $f_{0}\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 2}$. Set the hypothesis class as a deep ReQU neural network class ${\mathcal{F}}={\mathcal{N}}(L,S)$ with $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d}{d+4s}})$. Let $\widehat{f}_{{\mathcal{D}}}^{\lambda}$ be the deep Sobolev regressor defined in (11) with regularization parameter $\lambda>0$. Then it follows that
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}$ $\leq{\mathcal{O}}(\lambda^{2})+{\mathcal{O}}\Big{(}n^{-\frac{2s}{d+4s}}\Big{)},$
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{]}$ $\leq{\mathcal{O}}(\lambda)+{\mathcal{O}}\Big{(}\lambda^{-1}n^{-\frac{2s}{d+4s}}\Big{)}.$
Further, setting the regularization parameter $\lambda={\mathcal{O}}(n^{-\frac{s}{d+4s}})$ implies
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}$ $\leq{\mathcal{O}}\Big{(}n^{-\frac{2s}{d+4s}}\Big{)},$
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{]}$ $\leq{\mathcal{O}}\Big{(}n^{-\frac{s}{d+4s}}\Big{)}.$
Here we omit the logarithmic factors and the constant behind the big ${\mathcal{O}}$ notation is independent of $n$.
Theorem 5.4 provides theoretical guidance for the selection of the size of neural networks and the choice of regularization parameters. In comparison to the regularization parameter $\lambda={\mathcal{O}}(n^{-\frac{2s}{d+2s}})$ employed in Theorem 4.3, $\lambda={\mathcal{O}}(n^{-\frac{s}{d+4s}})$ utilized in Theorem 5.4 is much larger. The $L^{2}(\mu_{X})$-rate ${\mathcal{O}}(n^{-\frac{2s}{d+4s}})$ of the deep Sobolev regressor does not attain the minimax optimality. Furthermore, the convergence rate ${\mathcal{O}}(n^{-\frac{s}{d+4s}})$ for the derivatives is also slower than the minimax optimal rate ${\mathcal{O}}(n^{-\frac{2(s-1)}{d+2s}})$ derived in Stone (1982).
5.2 Analysis for Semi-Supervised Deep Sobolev Regressor
@@
There exists absolute positive constants $\{B_{1,k}\}_{k=1}^{d}$, such that $\sup_{x\in\Omega}|D_{k}f_{0}(x)|\leq B_{1,k}$ for $1\leq k\leq d$. Further, the first-order partial derivatives of functions in hypothesis class ${\mathcal{F}}$ are also bounded, i.e., $\sup_{x\in\Omega}|D_{k}f(x)|\leq B_{1,k}$ for each $1\leq k\leq d$ and $f\in{\mathcal{F}}$. Denote by $B_{1}^{2}:=\sum_{k=1}^{d}B_{1,k}^{2}$
@@
Lemma 5.5 (Oracle inequality).
Suppose Assumptions 1 to 6 hold. Let $\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}$ be the semi-supervised deep Sobolev regressor defined in (11). Then it follows that for each $\lambda>0$, $n\geq\operatorname{VCdim}({\mathcal{F}})$ and $m\geq\operatorname{VCdim}(\nabla{\mathcal{F}})$,
@@
$\lesssim\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m),$
@@
$\lesssim\tilde{\beta}\lambda+\lambda^{-1}\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\lambda^{-1}\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)+\lambda^{-1}\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m),$
where $\tilde{\beta}$ is a positive constant defined as $\tilde{\beta}=\beta+B_{1}^{2}$, the approximation error $\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)$ and the generalization error $\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)$ are defined as those in Lemma 5.2. The generalization error $\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m)$ corresponding to the regularization term are defined as
$\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m)=\inf_{\delta>0}\Big{\{}B_{1}^{2}\frac{\operatorname{VCdim}(\nabla{\mathcal{F}})}{m\delta}+B_{1}\delta\Big{\}}.$
In comparison to Lemma 5.2, the error bound has not undergone significant changes, and it has only been augmented by one additional generalization error associated with the regularization term. Further, this term vanishes as the number of unlabeled data increases.
@@
$\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m+n)=\inf_{\delta>0}\Big{\{}B_{1}^{2}\frac{\operatorname{VCdim}(\nabla{\mathcal{F}})}{(n+m)\delta}+B_{1}\delta\Big{\}}.$
@@
Theorem 5.6 (Convergence rates).
@@
Suppose Assumptions 1 to 6 hold. Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. Assume that $f_{0}\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 2}$. Set the hypothesis class as a deep ReQU neural network class ${\mathcal{F}}={\mathcal{N}}(L,S)$ with $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d}{d+4s}})$. Let $\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}$ be the regularized empirical risk minimizer defined as (9) with regularization parameter $\lambda={\mathcal{O}}(n^{-\frac{s}{d+4s}})$. Then it follows that
${\mathbb{E}}_{({\mathcal{D}},{\mathcal{S}})\sim\mu^{n}\times\nu_{X}^{m}}\Big{[}\|\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}$ $\leq{\mathcal{O}}\Big{(}n^{-\frac{2s}{d+4s}}\Big{)}+{\mathcal{O}}\Big{(}n^{\frac{d}{2(d+4s)}}m^{-\frac{1}{2}}\Big{)},$
${\mathbb{E}}_{({\mathcal{D}},{\mathcal{S}})\sim\mu^{n}\times\nu_{X}^{m}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{]}$ $\leq{\mathcal{O}}\Big{(}n^{-\frac{s}{d+4s}}\Big{)}+{\mathcal{O}}\Big{(}n^{\frac{d+2s}{2(d+4s)}}m^{-\frac{1}{2}}\Big{)},$
where we omit the logarithmic factors and the constant behind the big ${\mathcal{O}}$ notation is independent of $n$.
@@
As the number of unlabeled data $m$ approaches infinity, the convergence rate of the semi-supervised deep Sobolev regressor tends to the rate derived in Theorem 5.4.
6 Applications and Numerical Experiments

### v2

@@ V-A Analysis for Deep Sobolev Regressor
Lemma V.2 (Oracle inequality).
Suppose Assumptions 1 to 5 hold. Let $\widehat{f}_{{\mathcal{D}}}^{\lambda}$ be the deep Sobolev regressor defined as (14). Then it follows that for each $\lambda>0$ and each $n\geq\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))$,
$\lesssim\beta\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\varepsilon_{\operatorname{gen}}({\mathcal{F}},n),$
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{]}$
$\lesssim\beta\lambda+\lambda^{-1}\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)+\lambda^{-1}\varepsilon_{\operatorname{gen}}({\mathcal{F}},n),$
@@
$\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)$
$=\inf_{f\in{\mathcal{F}}}\Big{\{}\|f-f_{0}\|_{L^{2}(\mu_{X})}^{2}+\lambda\|\nabla(f-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{\}},$
$\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)$
$=\frac{B_{0}^{2}+\sigma^{2}}{\log^{-1}n}\inf_{\delta>0}\Big{\{}\Big{(}\frac{2\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))}{n}\Big{)}^{\frac{1}{2}}+\delta\Big{\}}.$
As discussed in Section IV[1, Chapter 21] has investigated an optimization problem similar to the deep Sobolev regressor. However, to the best of our knowledge, we are the first to demonstrate the oracle inequality for the gradient of estimator. The proof employs a similar technique as that of Lemma V.1. Specifically, the deep Sobolev regressor acts as the minimizer of (14), which implies that it satisfies a variational inequality derived from the first-order optimality condition [89, 90]. By utilizing standard techniques from statistical learning theory, we are able to derive the desired oracle inequality.
In simple terms, if we select an appropriate neural network class and have a sufficiently large number of labeled data pairs, we can make the approximation error and generalization error arbitrarily small. Consequently, the overall error is primarily determined by the regularization parameter $\lambda$. At this point, the error bound aligns with rates in Lemma V.1.
Recall the oracle inequality derived in Lemma IV.1, which requires the neural network to approximate the regression function while restricting its gradient norm. In contrast, the approximation term in Lemma V.2 necessitates the neural network to approximate both the regression function and its derivatives simultaneously. Thus, we now introduce the following approximation error bound in $H^{1}$-norm.
Lemma V.3 (Approximation in $H^{1}$-norm).
Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. Set the hypothesis class as a deep ReQU neural network ${\mathcal{F}}={\mathcal{N}}(L,W,S)$ with $L={\mathcal{O}}(\log N)$ and $S={\mathcal{O}}(N^{d})$. Then for each $\phi\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 2}$, there exists $f\in{\mathcal{F}}$ such that
@@
Theorem V.4 (Convergence rates).
Suppose Assumptions 1 to 5 hold. Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. Assume that $f_{0}\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 2}$. Set the hypothesis class as a deep ReQU neural network class ${\mathcal{F}}={\mathcal{N}}(L,W,S)$ with $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d}{d+4s}})$. Let $\widehat{f}_{{\mathcal{D}}}^{\lambda}$ be the deep Sobolev regressor defined in (14) with regularization parameter $\lambda>0$. Then it follows that
$\leq{\mathcal{O}}(\lambda^{2})+{\mathcal{O}}\Big{(}n^{-\frac{2s}{d+4s}}\log^{4}n\Big{)},$
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{]}$
$\leq{\mathcal{O}}(\lambda)+{\mathcal{O}}\Big{(}\lambda^{-1}n^{-\frac{2s}{d+4s}}\log^{4}n\Big{)}.$
Further, setting $\lambda={\mathcal{O}}(n^{-\frac{s}{d+4s}}\log^{2}n)$ implies
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0}\|_{L^{2}(\mu_{X})}^{2}\Big{]}$ $\leq{\mathcal{O}}\Big{(}n^{-\frac{2s}{d+4s}}\log^{4}n\Big{)},$
${\mathbb{E}}_{{\mathcal{D}}\sim\mu^{n}}\Big{[}\|\nabla(\widehat{f}_{{\mathcal{D}}}^{\lambda}-f_{0})\|_{L^{2}(\nu_{X})}^{2}\Big{]}$ $\leq{\mathcal{O}}\Big{(}n^{-\frac{s}{d+4s}}\log^{2}n\Big{)}.$
Here the constant behind the big ${\mathcal{O}}$ notation is independent of $n$.
Theorem V.4 provides theoretical guidance for the selection of the size of neural networks and the choice of regularization parameters. In comparison to the regularization parameter $\lambda={\mathcal{O}}(n^{-\frac{2s}{d+2s}}\log^{3}n)$ employed in Theorem IV.3, $\lambda={\mathcal{O}}(n^{-\frac{s}{d+4s}}\log^{2}n)$ utilized in Theorem V.4 is much larger. The $L^{2}(\mu_{X})$-rate ${\mathcal{O}}(n^{-\frac{2s}{d+4s}})$ of the deep Sobolev regressor does not attain the minimax optimality. Furthermore, the convergence rate ${\mathcal{O}}(n^{-\frac{s}{d+4s}}\log^{4}n)$ for the derivatives is also slower than the minimax optimal rate ${\mathcal{O}}(n^{-\frac{2(s-1)}{d+2s}})$ derived in [85].
V-B Analysis for Semi-Supervised Deep Sobolev Regressor
@@
There exists positive constants $\{B_{1,k}\}_{k=1}^{d}$, such that $\sup_{x\in\Omega}|D_{k}f_{0}(x)|\leq B_{1,k}$ for $1\leq k\leq d$. Further, the first-order partial derivatives of functions in hypothesis class ${\mathcal{F}}$ are also bounded, i.e., $\sup_{x\in\Omega}|D_{k}f(x)|\leq B_{1,k}$ for each $1\leq k\leq d$ and $f\in{\mathcal{F}}$. Denote by $B_{1}^{2}:=\sum_{k=1}^{d}B_{1,k}^{2}$
@@
Lemma V.5 (Oracle inequality).
Suppose Assumptions 1 to 6 hold. Let $\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}$ be the semi-supervised deep Sobolev regressor defined in (14). For each $\lambda>0$, $n\geq\log N(B_{0}\delta,{\mathcal{F}},L^{2}({\mathcal{D}}))$ and $m\geq\max_{1\leq k\leq d}\log N(B_{1,k}\delta,D_{k}{\mathcal{F}},L^{2}({\mathcal{S}}))$,
@@
$\lesssim\tilde{\beta}\lambda^{2}+\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)$
$\quad+\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)+\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m),$
@@
$\lesssim\tilde{\beta}\lambda+\lambda^{-1}\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)$
$\quad+\lambda^{-1}\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)+\lambda^{-1}\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m),$
where $\tilde{\beta}$ is a positive constant defined as $\tilde{\beta}=\beta+B_{1}^{2}$, the approximation error $\varepsilon_{\operatorname{app}}({\mathcal{F}},\lambda)$ and the generalization error $\varepsilon_{\operatorname{gen}}({\mathcal{F}},n)$ are defined as those in Lemma V.2. The generalization error $\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m)$ corresponding to the regularization term are defined as
$\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m)=B_{1}^{2}\inf_{\delta>0}\Big{\{}\max_{1\leq k\leq d}\frac{N(B_{1,k}\delta,D_{k}{\mathcal{F}},L^{2}({\mathcal{S}}))}{m}+\delta\Big{\}}.$
In comparison to Lemma V.2, the error bound has not undergone significant changes, and it has only been augmented by one additional generalization error associated with the regularization term. Further, this term vanishes as the number of unlabeled data increases.
@@
$\varepsilon_{\operatorname{gen}}^{\operatorname{reg}}(\nabla{\mathcal{F}},m+n)\\ =B_{1}^{2}\inf_{\delta>0}\Big{\{}\max_{1\leq k\leq d}\frac{\log N(B_{1,k}\delta,D_{k}{\mathcal{F}},L^{2}({\mathcal{S}}))}{m+n}+\delta\Big{\}}.$
@@
Theorem V.6 (Convergence rates).
@@
Suppose Assumptions 1 to 6 hold. Let $\Omega\subseteq K\subseteq{\mathbb{R}}^{d}$ be two bounded domain. Assume that $f_{0}\in C^{s}(K)$ with $s\in{\mathbb{N}}_{\geq 2}$. Set the hypothesis class as a deep ReQU neural network class ${\mathcal{F}}={\mathcal{N}}(L,W,S)$ with $L={\mathcal{O}}(\log n)$ and $S={\mathcal{O}}(n^{\frac{d}{d+4s}})$. Let $\widehat{f}_{{\mathcal{D}},{\mathcal{S}}}^{\lambda}$ be the regularized empirical risk minimizer defined as (12) with regularization parameter $\lambda={\mathcal{O}}(n^{-\frac{s}{d+4s}}\log^{2}n)$. Then it follows that
$\leq{\mathcal{O}}\Big{(}n^{-\frac{2s}{d+4s}}\log^{4}n\Big{)}+{\mathcal{O}}\Big{(}n^{\frac{d}{d+4s}}\log^{4}nm^{-1}\Big{)},$
$\leq{\mathcal{O}}\Big{(}n^{-\frac{s}{d+4s}}\log^{2}n\Big{)}+{\mathcal{O}}\Big{(}n^{\frac{d+s}{d+4s}}\log^{2}nm^{-1}\Big{)}.$
Here the constant behind the big ${\mathcal{O}}$ notation is independent of $n$.
@@
For the number of unlabeled data $m$ sufficiently large, the convergence rate of the semi-supervised deep Sobolev regressor tends to the rate derived in Theorem V.4.
VI Applications and Numerical Experiments